OpenStudy (anonymous):

\sqrt[3]{\left(\frac{1}{27}\right)^2}

OpenStudy (anonymous):

\[\sqrt[3]{\left(\frac{1}{27}\right)^2}~~?\]

OpenStudy (anonymous):

\[\sqrt[3]{\left(\frac{1}{27}\right)^2}\]